toni has a collection of dimes & quarters with a face value of $25.30. She has 27 more quarters than dimes. How many of each coin does she have?
Answer by oberobic(2304)

With money problems you have to keep track of the count of each type of coin as well as their value.
d = number of dimes
10d = value of the dimes in cents
q = number of quarters
25q = value of quarters in cents
$25.30 = 2530 cents (because you need to convert everything to cents)
10d + 25q = 2530
.
We also know she has 27 more quarters than dimes.
q = d + 27
.
Substituting q = d+27 into the total value equation, we have:
10d + 25(d+27) = 2530
10d + 25d + 675 = 2530
35d = 1855
d = 53
.
q = d + 27
q = 53 + 27
q = 80
.
Check the values
10d = 530
25q = 2000
10d + 25q = 2530
Checks.
.
So she has 53 dimes and 80 quarters.
.
Done.